Database last updated on Mon, 02 Feb 2026 10:10:52 +0000 by localhost
Instrument constant changes for the week before 2023307 (1 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 17 rows)
Measurements and tests summary.
_jday
_loc
_serial
_skc
dt
ed
fm
fz
hg
rs
sl
uv
307
RESOLUTE
#031
ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.186)
1
2
58
3
12
1
6
46
(0 sec)
Dead time (0 sec)
File name
Count
Serial
Expected (+/-5)
Measured
30-day history
B30723.031
1
#031
35
34